OpenStudy (tylermckinney16):

Calculate the cost of the trip: Gas: 400 miles divided by 30 miles per gallon times $2.75 per gallon Lodging: Cheap and free campgrounds with a tent. Stretching trip over five days totals $75 Food: Slightly cheaper from cooking meals at the campground. $350

OpenStudy (tylermckinney16):

@brittanyrolfe

OpenStudy (brittanyrolfe):

Okay so let's start with gas. \[\left( \frac{ 400 }{ 30 } \right)\left(2.75 \right)\] We are multiplying these together. So work inside the first parentheses and divide 400 by 30 (mile to the gallon) and then multiply that total by 2.75 (price of gas)

OpenStudy (brittanyrolfe):

What do you get as your answer for solving for the price of gas? @TylerMckinney16

OpenStudy (tylermckinney16):

36.6666666666 is what i got after doing what u said.

OpenStudy (brittanyrolfe):

Okay good. Lets round that to $36.67 since we are talking about money.

OpenStudy (brittanyrolfe):

Now we are going to add that to the other number present in the example since we are looking for the total cost. So $36.67 for gas + $75 for lodging +$350 for food

OpenStudy (brittanyrolfe):

Add those up and tell me what you get

OpenStudy (tylermckinney16):

$461.67

OpenStudy (brittanyrolfe):

That is your answer :)